Silicon of symbol SI is a metalloid used in fabrication of solar cells and transistors. It's prepared from pure sand. Silicon has 4 valence electrons and it's is located in row 3 in the periodic table. The number of nucleons in the silicon atom is 28.

1)Write the Lewis dot structure of the silicon atom.

2)write the electron configuration of silicon and justify.

3)Determine its atomic number.

4)Calculate the number of uncharged particles in the silicon atom.

5)determine the atomic number of the element just above the silicon atom in the periodic table.


1
Expert's answer
2021-03-02T01:03:26-0500

(i) Lewis dot structure of sillicon is-



(ii) Electronic configuration-"1s^2,2s^2,2p^6,3s^2,3p^2"

As It has 4 valence electron.


(iii) Atomic number of Sillicon is 14.


(iv) Number of uncharge particl=Number of Nucleons- Numberof protons

"=28-14=14"


(v) Atomic number of atom just above sillicon is =14-8=6